Zack is 2 years older than Jimmy. Last year Zack was 3 times as old as Jimmy. How is Zack now?

Answer by MILIAJEY(4)

LET JIMMY BE X YEARS
THEN TODAY ZACK 'S AGE Y= X+2
ONE YEAR BACK AGE OF ZACK (Y-1)= 3(X-1)
= 3X -3
Y=3X-3 +1
= 3X -2
BUT Y= X + 2
X+ 2 = 3X -2
4 = 3X-X
4 = 2X
2 = X
Y = 4
JIMMY IS 2 YRS OLD AND ZACK IS 4 YRS OLD.